# Group Therapy for Halos: Advancing Halo Mass Estimation for Galaxy Groups

### v1

@@ 3.1.2 MCMC Calibration
where $\alpha$, $\sigma_{lim}$, $n_{1}$, $\beta$, $R_{lim}$, and $n_{2}$ are free parameters determined by the MCMC calibration. The resulting posterior distributions for all model parameters are presented in Figure A8 in A, demonstrating that the parameters are well-constrained. The best-fitting parameters for the calibrated virial theorem are given in Table 1 for convenience.
@@
Notably, the calibrated $A_{\sigma}$ and $A_{R}$ coefficients decay to zero at the fitted velocity dispersion and group radius limits, reflecting that for larger multiplicity groups, the velocity dispersion and radius are well-defined and no further correction is required. Above these limits, the calibrated virial theorem (Equation 6) reduces to the standard virial theorem (Equation 3). This calibration procedure ensures that the virial theorem-based estimator is unbiased and robust across the full range of group properties, with well-quantified uncertainties and minimal model dependence.
To convert halo masses between different values of the Hubble parameter $H_{0}$ (or $h$), the following scaling should be applied:
@@
where $R_{lim}^{{}^{\prime}}$ is the group radius limit for a newly chosen $h^{\prime}$, and $h$ is the value used in this work ($h=0.7$). For example, converting to $h^{\prime}=0.67$ yields $R_{lim}^{{}^{\prime}}=0.369\cdot\frac{0.7}{0.67}=0.386$ Mpc. Alternatively, a factor of $\log_{10}(h^{\prime}/h)$ can be added to the calculated $M_{\mathrm{halo}}$ from Equation 6 when using the $R_{lim}$ in Table 1. The group radius for a given group should also be recalculated for the chosen $h$ value.
3.1.3 Accuracy, Uncertainty and Use Cases
We evaluate the accuracy and uncertainty of the calibrated virial-theorem-based halo mass estimator (Equation 6) using three independent SAMs: Shark, SAGE, and GAEA. The bottom panels of Figure 3 presents a quantitative comparison between the predicted halo mass from the calibrated virial theorem (Equation 6) and true halo masses for each model (calibrated to $z<0.1$ and $i<19.2$). The calibrated relation yields a high degree of consistency in both normalisation and slope across the models compared to the traditional virial theorem estimator (top panels), with only minor differences from the true halo mass, particularly at the low-mass end in the GAEA model.
@@ 3.1.3 Accuracy, Uncertainty and Use Cases
The accuracy of the calibrated virial theorem is characterised by the mean offset (mean $\Delta$) between the predicted and true halo masses, while the uncertainty is quantified by the mean of the $16^{\mathrm{th}}$ and $84^{\mathrm{th}}$ percentiles (mean $\sigma$) of the residuals. These metrics are shown in the bottom panels of Figure 3 as a function of the predicted halo mass for each simulation. For the Shark model, the mean $\Delta$ is 0.01-0.01 dex, with a mean $\sigma$ of $0.20$ dex, indicating negligible systematic bias and moderate scatter. The SAGE model yields a mean $\Delta$ of $0.12$ dex and a mean $\sigma$ of $0.10$ dex, while the GAEA model exhibits a mean $\Delta$ of 0.04-0.04 dex and a mean $\sigma$ of $0.13$ dex. These results demonstrate that the calibration procedure effectively removes systematic biases and reduces scatter across the full mass range and for all group multiplicities. The small variance in the calibrated relation between the different models reflects the minimal dependence of the virial theorem on the details of baryonic physics, as it is fundamentally anchored in gravitational dynamics. However, some residual differences are observed, particularly in the low-mass regime for the GAEA simulation.
@@
The robustness of the calibrated relation was further tested by varying the magnitude and redshift limits. When imposing a fainter WAVES-wide magnitude limit ($Z<21.2$), the free parameters changed by 1–20%, with the correction factor decreasing due to the deeper limiting magnitude. Extending the redshift range to $z<0.3$ resulted in larger changes of 5–40% in the free parameters, as stronger corrections were required as a function of redshift. These tests demonstrate that the calibrated relation presented here is specifically tailored to GAMA-like selection criteria (i.e., $i<19.2$) and to $z<0.1$. Users applying this relation to datasets with different selection functions or redshift limits should be aware of the potential systematic biases that may arise when when doing so and a recalibration of the relation for their specific survey parameters is suggested. Nevertheless, despite these limitations, the application of the calibrated relation to non-ideal datasets remains preferable to relying solely on the traditional virial theorem, which does not account for selection effects or redshift-dependent biases.
@@
In summary, the calibrated virial theorem is the least model-dependent method considered in this work, as it does not require knowledge of galaxy stellar masses or detailed prescriptions for baryonic processes amongst the models. This makes it ideally suited for applications where unbiased halo mass estimates are required, such as the construction of the halo mass function (HMF) in spectroscopic surveys or the derivation of cosmological parameters from group catalogues, provided that the relation is applied to the appropriate dataset.
3.2 Stellar-to-Halo Mass Relation
@@
Figure 4: Comparison of baryonic and halo mass relations across three SAMs: Shark, SAGE, and GAEA. The Shark sample has a fainter magnitude limit of $Z<21.2$ and a deeper redshift limit of $z<0.3$ compared to SAGE and GAEA ($i<19.2$ and $z<0.1$) Top Row: The relationship between the sum of the stellar masses of the three most massive group galaxies ($\log\Sigma M_{*,\mathrm{3}}$) and the true halo mass ($\log M_{\mathrm{halo}}$) is shown for each SAM, with the colour scale indicating the logarithm of the number of groups per bin. The black dashed line in each panel represents the MCMC-derived fit to the Shark data, optimised to minimise scatter in halo mass. This same fit is overlaid on the SAGE and GAEA panels to illustrate model dependence. Second Row: The residuals in halo mass ($\Delta\log M_{halo}=\log\Sigma M_{*,\,\mathrm{3}}-\log M_{\mathrm{halo}}$) are shown as a function of the summed stellar mass. Third Row: Estimated halo masses, derived by applying the Shark-calibrated relation to the summed stellar masses, are plotted against the true halo masses from each simulation. The black dashed line denotes the one-to-one relation. The close alignment of the points along this line in all models demonstrates that the Shark-based calibration provides robust halo mass estimates, with low bias and scatter, even when applied to independent SAMs. Bottom Row: The residuals in halo mass are shown as a function of the estimated halo mass. In both residual panels, the colour scale again indicates the logarithm of the number of groups per bin. These residual plots highlight the accuracy and systematic trends of the Shark-calibrated relation across the full range of halo and stellar masses, demonstrating low scatter and minimal bias, and further supporting the robustness of this approach for group-scale halo mass inference across diverse physical models.
@@
The connection between the stellar content of galaxies and their host dark matter haloes is a cornerstone of galaxy formation theory, providing a critical link between observable baryonic properties and the underlying dark matter distribution (Behroozi et al., 2010; Moster et al., 2013; Wechsler & Tinker, 2018). While abundance matching and halo occupation models have traditionally been used to infer this relationship on a statistical basis (e.g., Yang et al., 2003; Vale & Ostriker, 2004; Behroozi et al., 2010; Moster et al., 2013; Kravtsov et al., 2018), direct group-based approaches offer a powerful means to calibrate halo mass estimators for individual systems (e.g., Viola et al., 2015; Lim et al., 2021), particularly in the group regime where dynamical methods become less reliable at low multiplicity (Old et al., 2015; Muldrew et al., 2012).
@@ 3.2.1 Summed Stellar Mass Proxy
In this section, we establish an empirical relation between the sum of the stellar masses of the three most massive galaxies in a group ($\Sigma M_{*,\mathrm{3}}$) and the group halo mass ($M_{\mathrm{halo}}$). This approach is motivated by the expectation that the most massive group members should provide the most reliable baryonic tracers of the underlying halo mass. By summing the stellar mass of the three most massive galaxies, this method should mitigate stochasticity within the stellar-to-halo mass relation (SHMR) compared to using a single galaxy tracer.
The choice of three galaxies represents a balance between competing factors. Using only the most massive galaxy may introduce excessive scatter and result in a steeper high-mass slope in the SHMR, where small changes in stellar mass would produce disproportionately large changes in the estimated halo mass. Conversely, extending the sum to more galaxies may yield diminishing returns while adding noise from less reliable tracers. Summing over the three most massive galaxies should flatten the relation at the high-mass end and provide a more stable and physically motivated estimator.
Additionally, we assess the robustness of this relation by varying the selection criteria from the GAMA-like limit ($i<19.2$) to a fainter WAVES-wide limit ($Z<21.2$) and by extending the redshift range from $z<0.1$ to $z<0.3$, examining how the relation responds to different survey parameters.
3.2.2 MCMC Calibration and Functional Form
We utilise the same MCMC methodology as in Section 3.1.2. The likelihood function was constructed to minimise the scatter in $\log$ $M_{\mathrm{halo}}$ at fixed $\log\Sigma M_{\star,\,3}$; convergence was assessed via autocorrelation analysis. The resulting posterior distributions for the model parameters are shown in Figure B9, and further details of the likelihood function and fitting methodology are provided in B.
The adopted functional form for the relation is a double power-law:
@@ 3.2.2 MCMC Calibration and Functional Form
where $\Sigma M_{*,\mathrm{3}}$ is the sum of the stellar masses of the three most massive galaxies in the group in units of $M_{\odot}$, $A$ is the normalisation, $M_{A}$ is the characteristic stellar mass in units of $M_{\odot}$, and $\beta$ and $\gamma$ are the low- and high-mass slopes, respectively. The best-fitting parameters and their uncertainties, as determined from the MCMC analysis, are summarised in Table 2.
@@
where $M_{A}^{{}^{\prime}}$ is the characteristic stellar mass for a newly chosen $h^{{}^{\prime}}$, and h is the value used in this work ($h=0.7$). For example, converting to $h^{{}^{\prime}}=0.67$ yields $\log M_{A}^{{}^{\prime}}=\log(10^{10.483}\cdot\frac{0.7}{0.67})=10.502$. As the characteristic stellar mass is derived from simulations, there is only one factor of $h^{-1}$ rather than the typical $h^{-2}$ for observations. It is essential to ensure that both the stellar masses and the characteristic mass parameter $M_{A}$ are consistently defined with respect to the chosen value of $h$. Any conversion of the SHMR to a different Hubble parameter must be accompanied by a corresponding adjustment of these quantities.
3.2.3 Accuracy, Uncertainty and Use Cases
The accuracy and precision of the SHMR were assessed using the three independent SAMs: Shark, SAGE, and GAEA. As anticipated, using the combined stellar mass of the three most massive group members ($\Sigma M_{\star,~\mathrm{3}}$) proved to be the most effective approach. This method reduces the impact of stochasticity compared to using a single galaxy, while also avoiding the diminishing returns observed when incorporating more than three members. Importantly, extending the relation to include more than three galaxies would, by definition, exclude groups with fewer than four members from being appropriately sampled. Such groups would require a separate relation, introducing additional complexity and reducing the universality of the method. Relying solely on the most massive galaxy leads to a steep SHMR at the high-mass end, where small increases in stellar mass correspond to disproportionately large increases in halo mass. In contrast, the use of $\Sigma M_{\star,~\mathrm{3}}$ flattens this relation, producing a more gradual and stable increase in halo mass with stellar mass. This improvement enhances its suitability for observational studies, as uncertainties in stellar mass estimates translate to significantly smaller errors in the inferred halo mass.
The relation between $\Sigma M_{*,\mathrm{3}}$ and the true halo mass exhibits a high degree of consistency in overall form, but with notable differences in normalisation and slope between the models, as shown in Figure 4. These differences reflect the distinct implementations of baryonic physics, feedback, and merger histories in each SAM, which influence both stellar mass assembly and the mapping between baryonic and dark matter components (e.g., Somerville et al., 2015; Lacey et al., 2016; Lagos et al., 2018).
Testing the robustness of the relation across different selection criteria revealed that extending the redshift range to $z<0.3$ produced no significant changes, with fitted free parameters varying by less than 5%. The relation shown in Figure 4 is calibrated with Shark using the fainter WAVES-wide limit ($Z<21.2$) extended to $z<0.3$. This is what is shown within the Shark panels of Figure 4, whilst SAGE and GAEA both display a sample with ($i<19.2$) and $z<0.1$. Importantly, using the brighter GAMA-like limit ($i<19.2$) fails to capture the power-law slope at the low-mass end, making it unsuitable for calibrating this relation. The fainter $Z<21.2$ limit captures the low-mass regime, enabling a complete characterisation of the SHMR. Crucially, while the relation itself does not depend on the magnitude limit, the ability to define this relation does. This means the calibrated relation can be applied to any selection function with a brighter limit than $Z<21.2$, making it applicable to most existing group catalogues. Furthermore, since this method uses stellar mass rather than luminosity as the mass tracer, it can be applied to any group catalogue where reliable stellar mass measurements are available.
@@ 3.2.3 Accuracy, Uncertainty and Use Cases
As with the calibrated virial theorem in Section 3.2.3 the accuracy of the SHMR calibration is characterised by the mean $\Delta$ between the predicted and true halo masses, while the uncertainty is quantified by the mean $\sigma$. Both are presented in the second and bottom rows of Figure 4 as a function of both the sum of the primary three stellar masses ($\Sigma M_{*,\,\mathrm{3}}$) and the predicted halo mass ($M_{\mathrm{halo,\,\mathrm{3}}}$) for each simulation. For the Shark model, the mean $\Delta$ is $0.01$ dex as a function of the summed masses, with a mean $\sigma$ of $0.12$ dex; as a function of the predicted halo mass, the mean $\Delta$ is $0.02$ dex with a mean $\sigma$ of $0.14$ dex, indicating negligible systematic bias. Compared to the calibrated virial theorem in Section 3.1, the SHMR achieves similar accuracy but with approximately half the uncertainty in the Shark model.
@@
The SAGE and GAEA models yield mean $\Delta$ values of $0.21$ dex and $0.37$ dex, respectively, and mean $\sigma$ values of $0.10$ dex and $0.12$ dex, respectively, as a function of the summed stellar masses. As a function of the predicted halo mass, SAGE and GAEA yield mean $\Delta$ values of $0.13$ dex and $0.33$ dex, respectively, and mean $\sigma$ values of $0.11$ dex and $0.12$ dex, respectively. The differences in the SHMR between models are more pronounced than those seen for the calibrated virial theorem method. These discrepancies arise from the varying treatments of star formation efficiency, feedback, and merger-driven stellar mass growth in each SAM, as well as differences in the underlying dark matter merger trees and halo assembly histories (Mitchell et al., 2016; Fontanot et al., 2021). Such model dependencies are an inherent limitation of baryonic tracers, as the stellar mass content of haloes is sensitive to the adopted physical prescriptions and calibration strategies.
@@
It is important to note that, while the SHMR provides the most precise and lowest-scatter halo mass estimates, it is also the most model-dependent of the methods considered. The distribution of halo masses is tied to both the HMF and SMF of the input model—Shark. As a result, this halo mass estimator is not suitable for observationally constructing the HMF or for applications that aim to produce unbiased cosmological parameters from halo masses. For such applications, the calibrated virial theorem method is preferred, as its model dependence is negligible. Nevertheless, for studies focused on group-scale halo mass estimation, where the primary goal is to minimise residuals and uncertainties and to obtain a precise halo mass measurement, the SHMR is the optimal choice, provided that reliable stellar mass measurements are available.
@@
In summary, the calibrated SHMR offers a powerful and practical tool for estimating group halo masses, delivering the highest precision among the methods tested. Its application should be restricted to contexts where model dependence is not a limiting factor. The method’s versatility is demonstrated by its applicability to any group catalogue with stellar mass measurements and its robustness across different survey selection functions, provided they have magnitude limits brighter than $Z<21.2$. Critically, users must ensure that their stellar mass measurements are analogous to ProSpect-derived stellar masses (Robotham et al., 2020), as the Shark model used for this calibration is specifically tailored to ProSpect stellar masses. Failure to account for any systematic differences between stellar masses may introduce biases in the resulting halo mass estimates, particularly at the high-mass end.

### v3

@@ 4.1.2 Calibration of the Modified Virial Theorem
where $\alpha$, $\sigma_{lim}$, $n_{1}$, $\beta$, $R_{lim}$, and $n_{2}$ are free parameters determined by the MCMC sampling. The resulting posterior distributions for all model parameters are presented in Figure A9 in A, demonstrating that the parameters are well-constrained. The best-fitting parameters for the MVT are given in Table 1 for convenience.
@@
Notably, the calibrated $A_{\sigma}$ and $A_{R}$ coefficients decay to zero at the fitted velocity dispersion and group radius limits, reflecting that for larger multiplicity groups, the velocity dispersion and radius are well-defined and no further modification is required. Above these limits, the MVT (Equation 6) reduces to the virial theorem (Equation 3). This procedure ensures that the MVT estimator is unbiased and robust across the full range of group properties, with well-quantified uncertainties and minimal model dependence.
To convert halo masses between different values of the Hubble parameter $H_{0}$ (or Hubble constant $h$), the following scaling should be applied:
@@
where $R_{lim}^{{}^{\prime}}$ is the group radius limit for a newly chosen $h^{\prime}$, and $h$ is the value used in this work ($h=0.7$). For example, converting to $h^{\prime}=0.67$ yields $R_{lim}^{{}^{\prime}}=0.369\cdot\frac{0.7}{0.67}=0.386$ Mpc. Alternatively, a factor of $\log_{10}(h^{\prime}/h)$ can be added to the calculated $M_{\mathrm{halo}}$ from Equation 6 when using the $R_{lim}$ in Table 1. The calculated projected group radius for a given group should also be recalculated for the chosen $h$ value.
4.1.3 MVT - Accuracy, Uncertainty and Use Cases
@@
The MVT is calibrated to Shark, we then quantify the calibration accuracy and uncertainty on Shark and cross-validate the calibrated estimator to both SAGE and GAEA. The bottom panels of Figure 4 present a quantitative comparison between the predicted halo mass from the MVT (Equation 6) and true halo masses for each model (calibrated to $z<0.1$ and $i<19.2$ mag). The accuracy of the MVT is characterised by the mean offset (mean $\Delta$) between the predicted and true halo masses, while the uncertainty is quantified by the mean of the $16^{\mathrm{th}}$ and $84^{\mathrm{th}}$ percentiles (mean $\sigma$) of the residuals. These metrics are shown in the bottom panels of Figure 4 as a function of the predicted halo mass for each simulation. On the calibration model (Shark), the mean $\Delta$ is 0.01-0.01 dex, with a mean $\sigma$ of $0.20$ dex, indicating negligible systematic bias and moderate scatter. On the validation models; SAGE yields a mean $\Delta$ of $0.12$ dex and a mean $\sigma$ of $0.10$ dex, while the GAEA model exhibits a mean $\Delta$ of $0.07$ dex and a mean $\sigma$ of $0.13$ dex. These results demonstrate that the calibration procedure effectively removes systematic biases and reduces scatter across the full mass range and for all group multiplicities. The calibrated relation yields a high degree of consistency in both normalisation and slope across the models compared to the traditional virial theorem estimator (top panels), with only minor differences from the true halo mass, particularly at the low-mass end in the GAEA model. The small variance in the calibrated relation between the different models reflects the minimal dependence of the virial theorem on the details of baryonic physics, as it is fundamentally anchored in gravitational dynamics. However, some residual differences are observed, particularly in the low-mass regime for the GAEA simulation.
@@
The robustness of the calibrated relation was further tested by varying the magnitude and redshift limits. When imposing a fainter WAVES-wide magnitude limit ($Z<21.2$), the free parameters changed by 1–20%, with the correction factor decreasing due to the deeper limiting magnitude. Extending the redshift range to $z<0.3$ resulted in larger changes of 5–40% in the free parameters, as stronger corrections were required as a function of redshift. These tests demonstrate that the calibrated relation presented here is specifically tailored to GAMA-like selection criteria (i.e., $i<19.2$) and to $z<0.1$. Users applying this relation to datasets with different selection functions or redshift limits should be aware of the potential systematic biases that may arise when doing so and a recalibration of the relation for their specific survey parameters is suggested. Nevertheless, despite these limitations, the application of the calibrated relation to non-ideal datasets remains preferable to relying solely on the traditional virial theorem, which does not account for selection effects or redshift-dependent biases.
@@
In summary, the MVT is the least model-dependent method considered in this work, as it does not require knowledge of galaxy properties or detailed prescriptions for baryonic processes amongst the models. This makes it ideally suited for applications where unbiased halo mass estimates are required, such as the construction of the halo mass function (HMF) in spectroscopic surveys or the derivation of cosmological parameters from group catalogues, provided that the relation is applied to the appropriate dataset.
4.2 The summed Stellar-to-Halo Mass Relation
@@
Figure 5: Comparison of baryonic and halo mass relations across three SAMs: Shark, SAGE, and GAEA. The Shark sample has a fainter magnitude limit of $Z<21.2$ and a deeper redshift limit of $z<0.3$ compared to SAGE and GAEA ($i<19.2$ and $z<0.1$) Top Row: The sSHMR using the summed stellar mass of the three most massive group galaxies ($\log\Sigma M_{*,\mathrm{3}}$) and the true halo mass ($\log M_{\mathrm{halo}}$) is shown for each SAM, with the colour scale indicating the logarithm of the number of groups per bin. The black dashed line in each panel represents the MCMC-derived fit to the Shark data, optimised to minimise scatter in halo mass. This same fit is overlaid on the SAGE and GAEA panels to illustrate model dependence. Second Row: The residuals in halo mass ($\Delta\log M_{halo}=\log\Sigma M_{*,\,\mathrm{3}}-\log M_{\mathrm{halo}}$) is shown as a function of the summed stellar mass. Third Row: Estimated halo masses, derived by applying the Shark-calibrated sSHMR, are plotted against the true halo masses from each simulation. The black dashed line denotes the one-to-one relation. The close alignment of the points along this line in all models demonstrates that the Shark-based calibration provides robust halo mass estimates, with low bias and scatter, even when applied to independent SAMs. Bottom Row: The residuals in halo mass are shown as a function of the estimated halo mass. In both residual panels, the colour scale again indicates the logarithm of the number of groups per bin.
@@
The connection between the stellar content of galaxies and their host dark matter haloes is a cornerstone of galaxy formation theory, providing a critical link between observable baryonic properties and the underlying dark matter distribution (Behroozi2010; Moster2013; Wechsler2018). While abundance matching and halo occupation models have traditionally been used to infer this relationship on a statistical basis (e.g., Yang2003; Vale2004; Behroozi2010; Moster2013; Kravtsov2018), direct group-based approaches offer a powerful means to calibrate halo mass estimators for individual systems (e.g., Viola2015; Lim2021), particularly in the group regime where dynamical methods become less reliable at low multiplicity (Old2015; Muldrew2012).
@@ 4.2.1 Summed Stellar Mass Proxy
In this section, we establish an empirical sSHMR between the summed stellar masses of the three most massive galaxies in a group ($\Sigma M_{*,\mathrm{3}}$) and the group halo mass ($M_{\mathrm{halo}}$). This approach is motivated by the expectation that the most massive group members should provide the most reliable baryonic tracers of the underlying halo mass. By summing the stellar mass of the three most massive galaxies, this method should mitigate stochasticity compared to a typical single galaxy tracer via a stellar-to-halo mass relation (SHMR). The choice of three galaxies represents a balance between competing factors, using only the most massive galaxy may introduce excessive scatter and result in a steeper high-mass slope in the SHMR, where small changes in stellar mass would produce disproportionately large changes in the estimated halo mass. Conversely, extending the sum to more galaxies may yield diminishing returns while adding noise from less reliable tracers. Summing over the three most massive galaxies should flatten the relation at the high-mass end and provide a more stable and physically motivated estimator. Additionally, we assess the robustness of this relation by varying the selection criteria from the GAMA-like limit ($i<19.2$) to a fainter WAVES-wide limit ($Z<21.2$) and by extending the redshift range from $z<0.1$ to $z<0.3$, examining how the relation responds to different survey parameters.
4.2.2 sSHMR Calibration and Functional Form
We use the same MCMC methodology as in Section 4.1.2. The likelihood function was constructed to minimise the scatter in $\log$ $M_{\mathrm{halo}}$ at fixed $\log\Sigma M_{\star,\,3}$; convergence was assessed via autocorrelation analysis. The resulting posterior distributions for the model parameters are shown in Figure B10, and further details of the likelihood function, the fitting methodology and the uncertainty of the fitted parameters are provided in B.
The adopted functional form for the sSHMR is a double power-law in the form:
@@
where $\Sigma M_{*,\mathrm{3}}$ is the sum of the stellar masses of the three most massive galaxies in the group in units of $M_{\odot}$, $A$ is the normalisation, $M_{A}$ is the characteristic stellar mass in units of $M_{\odot}$, and $\beta$ and $\gamma$ are the low- and high-mass slopes, respectively. The best-fitting parameters from the MCMC sampling are summarised in Table 2.
@@
where $M_{A}^{{}^{\prime}}$ is the characteristic stellar mass for a newly chosen $h^{{}^{\prime}}$, and h is the value used in this work ($h=0.7$). For example, converting to $h^{{}^{\prime}}=0.67$ yields $\log M_{A}^{{}^{\prime}}=\log(10^{10.483}\cdot\frac{0.7}{0.67})=10.502$. As the characteristic stellar mass is derived from simulations, there is only one factor of $h^{-1}$ rather than the typical $h^{-2}$ for observations. It is essential to ensure that both the stellar masses used in the sSHMR and the characteristic mass parameter $M_{A}$ are consistently defined with respect to the chosen value of $h$. Any conversion of the sSHMR to a different Hubble parameter must be accompanied by a corresponding adjustment of these quantities.
4.2.3 sSHMR - Accuracy, Uncertainty and Use Cases
@@
Following the same methods as in Section 4.1.3, the sSHMR is calibrated to Shark where we test the accuracy and uncertainty of this calibration, then we-cross validate this with SAGE and GAEA as independent tests. The accuracy of the sSHMR calibration is characterised by the mean $\Delta$ between the predicted and true halo masses, while the uncertainty is quantified by the mean $\sigma$. Both are presented in the second and last rows of Figure 5 as a function of both the sum of the primary three stellar masses ($\Sigma M_{\star,\,3}$) and the predicted halo mass ($M_{\mathrm{halo,\,3}}$) for each model. On the calibration model (Shark), the mean $\Delta$ is $0.01$ dex as a function of the summed masses, with a mean $\sigma$ of $0.12$ dex; as a function of the predicted halo mass, the mean $\Delta$ is $0.02$ dex with a mean $\sigma$ of $0.14$ dex, indicating negligible systematic bias. Compared to the MVT in Section 4.1.3, the sSHMR achieves similar accuracy but with approximately half the uncertainty in the Shark model.
@@
Through cross-validation the SAGE and GAEA models yield mean $\Delta$ values of $0.21$ dex and $0.37$ dex, respectively, and mean $\sigma$ values of $0.10$ dex and $0.12$ dex, respectively, as a function of the summed stellar masses. As a function of the predicted halo mass, SAGE and GAEA yield mean $\Delta$ values of $0.13$ dex and $0.33$ dex, respectively, and mean $\sigma$ values of $0.11$ dex and $0.12$ dex, respectively. The differences in the sSHMR between models are more pronounced than those seen for the MVT. These discrepancies arise from the varying treatments of star formation efficiency, feedback, and merger-driven stellar mass growth in each SAM, as well as differences in the underlying dark matter merger trees and halo assembly histories. Such model dependencies are an inherent limitation of baryonic tracers, as the stellar mass content of halos is sensitive to the adopted physical prescriptions and calibration strategies.
As hypothesised, using a sSHMR proved to be the most effective approach rather than a typical SHMR. This method reduces the impact of stochasticity compared to using a single galaxy tracer, while also avoiding the diminishing returns observed when incorporating more than three members. Importantly, extending the relation to include more than three galaxies would, by definition, exclude groups with fewer than four members from being appropriately sampled. Such groups would require a separate relation, introducing additional complexity and reducing the universality of the method. Relying solely on the most massive galaxy leads to a steep slope at the high-mass end, where small increases in stellar mass correspond to disproportionately large increases in halo mass. In contrast, the use of $\Sigma M_{\star,\,3}$ flattens this relation, producing a more gradual and stable increase in halo mass with stellar mass. This improvement enhances its suitability for observational studies, as uncertainties in stellar mass estimates translate to significantly smaller errors in the inferred halo mass. The relation between $\Sigma M_{\star,\,3}$ and the true halo mass exhibits a high degree of consistency in overall form, but with notable differences in normalisation and slope between the models, as shown in Figure 5. These differences reflect the distinct implementations of baryonic physics, feedback, and merger histories in each SAM, which influence both stellar mass assembly and the mapping between baryonic and dark matter components.
Testing the robustness of the relation across different selection criteria demonstrated that extending the redshift range to $z<0.3$ produced no significant changes, with the fitted free parameters varying by less than 5%. The relation shown in Figure 5 is calibrated on Shark using the fainter WAVES-wide limit ($Z<21.2$) extended to $z<0.3$. This is what is shown within the Shark panels of Figure 5, whilst SAGE and GAEA both display a sample with ($i<19.2$) and $z<0.1$. Importantly, using the brighter GAMA-like limit ($i<19.2$) fails to capture the power-law slope at the low-mass end, making it unsuitable for calibrating this relation. The fainter $Z<21.2$ limit captures the low-mass regime, enabling a complete characterisation of the sSHMR. Crucially, while the relation itself does not depend on the magnitude limit, the ability to define this relation does. This means the calibrated relation can be applied to any selection function with a brighter limit than $Z<21.2$, making it applicable to most existing group catalogues. Furthermore, since this method uses stellar mass rather than luminosity as the mass tracer, it can be applied to any group catalogue where reliable stellar mass measurements are available.
@@
It is important to note that, while the sSHMR provides the most precise and lowest-scatter halo mass estimates, it is also the most model-dependent of the methods considered. The distribution of halo masses is tied to both the HMF and SMF of the input model—Shark. As a result, this halo mass estimator is not suitable for observationally constructing the HMF or for applications that aim to produce unbiased cosmological parameters from halo masses. For such applications, the MVT method is preferred, as its model dependence is negligible. Nevertheless, for studies focused on group-scale halo mass estimation, where the primary goal is to minimise residuals and uncertainties and to obtain a precise halo mass measurement, the sSHMR is the optimal choice, provided that reliable stellar mass measurements are available.
@@
In summary, the calibrated sSHMR offers a powerful and practical tool for estimating group halo masses, delivering the highest precision among the methods tested. Its application should be restricted to contexts where model dependence is not a limiting factor. The method’s versatility is demonstrated by its applicability to any group catalogue with stellar mass measurements and its robustness across different survey selection functions, provided they have magnitude limits brighter than $Z<21.2$. Critically, users must ensure that their stellar mass measurements are analogous to ProSpect-derived stellar masses, as the Shark model used for this calibration is specifically tailored to ProSpect stellar masses. Failure to account for any systematic differences between stellar masses may introduce biases in the resulting halo mass estimates, particularly at the high-mass end.